OpenStudy (anonymous):

the population of the earth is approximately 6.1 billion people and is growing at an annual rate of 1.4%. Assuming a Malthusian growth mode, find the world population in 29 years. Please round the answer to the nearest tenth of a billion.

OpenStudy (perl):

I think we can use this equation: $$ \Large y =6.1~\left(1 + \frac{1.4}{100} \right)^t $$

OpenStudy (anonymous):

thank you so much for helping me :)

OpenStudy (perl):

the "malthusian model' seems a little ambiguous, but the two models come out about the same $$\Large y= 6.1 e ^{0.014 t } $$

OpenStudy (perl):

actually it does make a difference, since it says round to the tenth place using the equation with e, i get 9.2 billion rounded to nearest tenth place

OpenStudy (anonymous):

thats what I got as well

OpenStudy (perl):

the two models come out differently, if you round out to the nearest tenth. i would go with the latter model, using e

OpenStudy (perl):

$$\Large { f(t) =6.1~\left(1 + .014 \right)^{~t} \\\Large f(29) =6.1~\left(1 + .014 \right)^{~29} = 9.12915 \\~\\\Large g(t)= 6.1 e ^{0.014 t } \\ g(29) = 6.1 e ^{0.014 \times 29 } = 9.154895 }$$